# Real-Time LiDAR Point Cloud Densification for Low-Latency Spatial Data Transmission

### v1

@@ Abstract
To realize low-latency spatial transmission system for immersive telepresence, there are two major problems: capturing dynamic 3D scene densely and processing them in real time. LiDAR sensors capture 3D in real time, but produce sparce point clouds. Therefore, this paper presents a high-speed LiDAR point cloud densification method to generate dense 3D scene with minimal latency, addressing the need for on-the-fly depth completion while maintaining real-time performance. Our approach combines multiple LiDAR inputs with high-resolution color images and applies a joint bilateral filtering strategy implemented through a convolutional neural network architecture. Experiments demonstrate that the proposed method produces dense depth maps at full HD resolution in real time (30 fps), which is over 15x faster than a recent training-based depth completion approach [BPNet]. The resulting dense point clouds exhibit accurate geometry without multi-view inconsistencies or ghosting artifacts.
@@ 1 Introduction
In recent years, many projects have explored capturing the entirety of live events (e.g. music concerts or sports) in both time and space and transmitting that data to remote locations to provide viewers with a sense of immersion as if they were physically present [holoportation, dou2016fusion4d]. We refer to a system that realizes such a service as a spatial transmission system. A spatial transmission system needs to capture a 3D space in real time, transmit it with low latency, and present it to users as a 3D representation. Among these requirements, the particularly challenging issue is the real-time capture of a dense 3D space.
Volumetric video capture [volumetric] can reconstruct 3D information from numerous camera feeds and enable free-viewpoint rendering, but it requires several seconds of processing and thus cannot achieve the low latency needed for interactive experiences. Moreover, they focus on only human’s behavior, not space. Alternatively, using 3D sensors like LiDAR allows real-time 3D capture; however, LiDAR data is much sparser than camera images. Achieving an immersive level of detail with LiDAR alone would require a very large number of sensors, which is impractical.
@@ 3 Related Work
Reconstructing a dense depth map from sparse depth data with the help of a high-resolution RGB image is called guided depth completion, and has been extensively studied [DepthCompSurvey2022, RGBGuidedSurvey2024]. For example, DAGF [DAGF] employs a guided filter-based CNN model that uses the RGB image as guidance to upsample a low-resolution depth image. This approach works well when the sparse depth samples lie on a regular grid. However, in our system the LiDAR’s viewpoint is different from the camera’s, so the LiDAR depth points appear at random positions when projected onto the camera image. Methods that assume grid-structured depth input, like DAGF, cannot be directly applied in this scenario.
Recently, many approaches try to densify randomly sparse depth with RGB image as guidance [RGBGuidedSurvey2024], and some of them have achieved real-time processing [BPNet, bai2020depthnet, long2024sparsedc]. BP-Net [BPNet] builds a CNN model that propagates the sparse depth information to neighboring pixels, guided by the RGB image, to fill in missing depths. It achieves high accuracy and is faster than many prior methods. However, BP-Net still requires about 20 ms to process a 320×240 image, which is insufficient for our goal of 30 fps at Full HD resolution. Moreover, most of existing methods rely on supervised learning with high-density depth data for training. In practice, obtaining such ground-truth dense depth for training is difficult, and methods trained on one domain (e.g. indoor datasets like NYU Depth [NYUDepthV2] ) may not generalize robustly to different environments. This need for extensive training data and the potential domain gap pose additional challenges for deploying these methods in a live system.
@@ 4 Proposed Method
Our proposed method focuses on processing high-resolution images efficiently, and is built around a fast, accurate depth densification approach based on bilateral filtering [BF]. Bilateral filtering (BF) is a classical edge-preserving smoothing technique originally used for image denoising, which combines color and spatial weighting. A variant called joint bilateral filtering (JBF) [JBF] uses a separate guidance image to guide the filtering. We leverage this concept for depth completion: the idea is to smooth and interpolate the sparse depth values in a way that respects edges in the high-resolution color image. We implement joint bilateral filtering in a CNN-inspired manner to fully exploit GPU parallelism and allow the filter parameters to be trainable. By formulating the filtering operations as convolution-like layers, our method can be executed rapidly using deep learning libraries (e.g. PyTorch [PyTorch]), and if ground-truth dense depth is available, parameters such as the Gaussian variances can be optimized via backpropagation. Importantly, our approach does not require pre-training on dense depth data. The filter parameters can be set empirically and still produce good results, with the option to fine-tune them if data is available.
@@ 4.2 Depth Densification via Bilateral Filtering
Our bilateral filter-based depth completion can be implemented very efficiently on a GPU. We have written it using PyTorch [PyTorch], taking advantage of optimized tensor operations. As a result, the entire filtering can run in real-time even on high-resolution images.
@@ 5 Experimental Results
Fig. 3 shows an example of the obtained results. The figure compares the point cloud before densification, the after densification result using our method, and the result produced by BP-Net [BPNet]. For the BP-Net result, we used a publicly available implementation with a model pre-trained on the NYU Depth v2 indoor dataset [NYUDepthV2]. From the figure, it is clear that our method successfully produces a much denser 3D point cloud of the person and the environment. The reconstructed surfaces using our approach appear complete and well-aligned across the multiple viewpoints. In contrast, BP-Net does increase the density of the point cloud to some extent and the depth maps from each view appear smooth; however, we observed that BP-Net tended to introduce a slight depth offset for the entire subject. When the depth outputs from different cameras are merged, these per-view depth misalignments cause the geometry to appear as ghosted double surfaces in the combined point cloud. Our method, despite its simplicity, yields depth maps that are properly aligned across all viewpoints, resulting in a clean integrated 3D reconstruction. The visual quality of our dense point cloud is on par with, or better than, the learning-based BP-Net in this multi-view scenario, while using a much simpler filtering model.

### v2

@@ Abstract
To realize low-latency spatial transmission system for immersive telepresence, there are two major problems: capturing dynamic 3D scene densely and processing them in real time. LiDAR sensors capture 3D in real time, but produce sparce point clouds. Therefore, this paper presents a high-speed LiDAR point cloud densification method to generate dense 3D scene with minimal latency, addressing the need for on-the-fly depth completion while maintaining real-time performance. Our approach combines multiple LiDAR inputs with high-resolution color images and applies a joint bilateral filtering strategy implemented through a convolutional neural network architecture. Experiments demonstrate that the proposed method produces dense depth maps at full HD resolution in real time (30 fps), which is over 15x faster than a recent training-based depth completion approach [10]. The resulting dense point clouds exhibit accurate geometry without multi-view inconsistencies or ghosting artifacts.
@@ 1 Introduction
In recent years, many projects have explored capturing the entirety of live events (e.g. music concerts or sports) in both time and space and transmitting that data to remote locations to provide viewers with a sense of immersion as if they were physically present [5, 2]. We refer to a system that realizes such a service as a spatial transmission system. A spatial transmission system needs to capture a 3D space in real time, transmit it with low latency, and present it to users as a 3D representation. Among these requirements, the particularly challenging issue is the real-time capture of a dense 3D space.
Volumetric video capture [8] can reconstruct 3D information from numerous camera feeds and enable free-viewpoint rendering, but it requires several seconds of processing and thus cannot achieve the low latency needed for interactive experiences. Moreover, they focus on only human’s behavior, not space. Alternatively, using 3D sensors like LiDAR allows real-time 3D capture; however, LiDAR data is much sparser than camera images. Achieving an immersive level of detail with LiDAR alone would require a very large number of sensors, which is impractical.
@@ 3 Related Work
Reconstructing a dense depth map from sparse depth data with the help of a high-resolution RGB image is called guided depth completion, and has been extensively studied [3, 7]. For example, DAGF [13] employs a guided filter-based CNN model that uses the RGB image as guidance to upsample a low-resolution depth image. This approach works well when the sparse depth samples lie on a regular grid. However, in our system the LiDAR’s viewpoint is different from the camera’s, so the LiDAR depth points appear at random positions when projected onto the camera image. Methods that assume grid-structured depth input, like DAGF, cannot be directly applied in this scenario.
Recently, many approaches try to densify randomly sparse depth with RGB image as guidance [7], and some of them have achieved real-time processing [10, 1, 4]. BP-Net [10] builds a CNN model that propagates the sparse depth information to neighboring pixels, guided by the RGB image, to fill in missing depths. It achieves high accuracy and is faster than many prior methods. However, BP-Net still requires about 20 ms to process a 320×240 image, which is insufficient for our goal of 30 fps at Full HD resolution. Moreover, most of existing methods rely on supervised learning with high-density depth data for training. In practice, obtaining such ground-truth dense depth for training is difficult, and methods trained on one domain (e.g. indoor datasets like NYU Depth [9] ) may not generalize robustly to different environments. This need for extensive training data and the potential domain gap pose additional challenges for deploying these methods in a live system.
@@ 4 Proposed Method
Our proposed method focuses on processing high-resolution images efficiently, and is built around a fast, accurate depth densification approach based on bilateral filtering [11]. Bilateral filtering (BF) is a classical edge-preserving smoothing technique originally used for image denoising, which combines color and spatial weighting. A variant called joint bilateral filtering (JBF) [12] uses a separate guidance image to guide the filtering. We leverage this concept for depth completion: the idea is to smooth and interpolate the sparse depth values in a way that respects edges in the high-resolution color image. We implement joint bilateral filtering in a CNN-inspired manner to fully exploit GPU parallelism and allow the filter parameters to be trainable. By formulating the filtering operations as convolution-like layers, our method can be executed rapidly using deep learning libraries (e.g. PyTorch [6]), and if ground-truth dense depth is available, parameters such as the Gaussian variances can be optimized via backpropagation. Importantly, our approach does not require pre-training on dense depth data. The filter parameters can be set empirically and still produce good results, with the option to fine-tune them if data is available.
@@ 4.2 Depth Densification via Bilateral Filtering
Our bilateral filter-based depth completion can be implemented very efficiently on a GPU. We have written it using PyTorch [6], taking advantage of optimized tensor operations. As a result, the entire filtering can run in real-time even on high-resolution images.
@@ 5 Experimental Results
Fig. 3 shows an example of the obtained results. The figure compares the point cloud before densification, the after densification result using our method, and the result produced by BP-Net [10]. For the BP-Net result, we used a publicly available implementation with a model pre-trained on the NYU Depth v2 indoor dataset [9]. From the figure, it is clear that our method successfully produces a much denser 3D point cloud of the person and the environment. The reconstructed surfaces using our approach appear complete and well-aligned across the multiple viewpoints. In contrast, BP-Net does increase the density of the point cloud to some extent and the depth maps from each view appear smooth; however, we observed that BP-Net tended to introduce a slight depth offset for the entire subject. When the depth outputs from different cameras are merged, these per-view depth misalignments cause the geometry to appear as ghosted double surfaces in the combined point cloud. Our method, despite its simplicity, yields depth maps that are properly aligned across all viewpoints, resulting in a clean integrated 3D reconstruction. The visual quality of our dense point cloud is on par with, or better than, the learning-based BP-Net in this multi-view scenario, while using a much simpler filtering model.